# Study of mass outflows from magnetized accretion disks around rotating black holes with thermal conduction

### v1

@@ IV Astrophysical Implications
Further, X-ray luminosity ($L_{\rm x}$) for the same sample of AGNs is calculated in $0.2-12$ keV energy range from XMM--Newton data Rosen et al. (2016), which spans in the range of $10^{40.99}-10^{45.02}$ $\rm erg\,s^{-1}$. Using $L_{\rm x}$, we estimate the accretion power as $\dot{M}c^{2}=L_{\rm x}/\eta_{\rm acc}$, where $\eta_{\rm acc}$ denotes the accretion efficiency factor. Since the spin measurements of supermassive black holes (SMBHs) in these AGNs are not well constrained, we adopt a maximum efficiency of $\eta_{\rm acc}=0.3$ in this study Thorne (1974). Thereafter, we employ the accretion-ejection model formalism to calculate the maximum kinetic jet power Aktar et al. (2015); Nandi et al. (2018); Jana and Das (2024) using maximum mass outflow rate $R^{\rm max}_{\dot{m}}$ as
@@ V SUMMARY and CONCLUSION
In this study, we examine the effect of thermal conduction on the mass loss from a relativistic, advective, viscous, and magnetized accretion disk around a rotating black hole, for the first time to the best of our knowledge. In doing so, we assume the accreting plasma is collisionless and adopt the saturated form of thermal conduction Cowie and McKee (1977). In addition, we model the accretion disk as being threaded with a dominant toroidal magnetic field component Oda et al. (2007); Sarkar and Das (2016); Sarkar et al. (2018); Jana and Das (2024). To simplify the general relativistic complexities, we adopt a recently developed effective potential Dihingia et al. (2018) that satisfactorily mimics the space-time around a rotating black hole. Furthermore, we employ the relativistic equation of state (REoS) Chattopadhyay and Ryu (2009) to determine the thermodynamic quantities. With these assumptions, we solve the coupled governing equations for both inflow and outflow self-consistently in terms of the inflow parameters, namely energy, angular momentum, plasma-$\beta$, and the conduction parameter. We summarize the key findings of this study below:
@@
  • We compute the maximum mass outflow rate $R^{\rm max}_{\dot{\rm m}}$ as a function of $\Upsilon_{\rm s}$ for two extreme black hole spin parameters, $a_{\rm k}=0.0$ and $0.99$ (Fig. 4). Our results reveal that rapidly rotating black holes expel more mass from the magnetized disk in the form of outflows compared to the weakly rotating black holes. More precisely, when thermal conduction is active within the disk, $R_{\dot{\rm m}}^{\rm max}$ reaches up to $24\%$ for $a_{\rm k}=0.99$, whereas it remains below $22\%$ for $a_{\rm k}=0.0$ (see Fig. 4).
@@ Acknowledgements
$R_{\rm v}$ $=v+T_{\rm v}/h,\quad R_{\Theta}=T_{\Theta}/h,\quad R_{\lambda}=T_{\lambda}/h,$
$R_{\beta}$ $=T_{\beta}/h,\quad{\rm and}\quad R_{0}=T_{0}/h+2t\Theta/(\beta x)+d\Psi_{\rm e}^{\rm eff}/dx,$
@@
$E_{\rm v}$ $=t\Theta+\frac{5\Upsilon(t\Theta)^{3/2}}{v},$
$\quad E_{\Theta}$ $=(vt/2+Nvt)-5\Upsilon t^{3/2}\sqrt{\Theta},$
@@
$+\frac{5\Upsilon(t\Theta)^{3/2}}{2}\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial\lambda},$
$E_{\beta}$ $=-\frac{vt\Theta}{2(1+\beta)\beta}-\frac{5\Upsilon(t\Theta)^{3/2}}{2\beta(1+\beta)},$
@@
$+5\Upsilon(t\Theta)^{3/2}\Big{(}-\frac{1}{x}+\frac{1}{2\Delta}\frac{d\Delta}{dx}+\frac{1}{2\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}\Big{)}.$
@@
$B_{0}$ $=\frac{1}{4}\Big{(}\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}-\frac{1}{\Delta}\frac{d\Delta}{dx})\Big{)}+\frac{\zeta}{x}.$
@@
Figure 6: Variation of (a) Mach number ($M=v/C_{\rm s}$) and (b) energy ($\mathcal{E}$) of the inflowing matter with radial distance $(x)$ for $\Upsilon_{\rm s}$ = $0.0$ (green), $0.0025$ (blue) and $0.005$ (red). Here, we choose $a_{\rm k}=0$, $\alpha_{\rm B}=0.005$, and fix the input parameters at $x_{\rm edge}=10^{3}$ as $\lambda_{\rm edge}=3.37$, $\mathcal{E}_{\rm edge}=4\times 10^{-4}$, and $\beta_{\rm edge}=10^{4}$. See the text for details.
Thermal conduction plays an important role in transferring heat energy within the accretion disk. Hence, it is worthy to examine the effect thermal conduction on the local energy of inflowing matter. In doing so, we choose $a_{\rm k}=0$, and $\alpha_{\rm B}=0.005$, and inject matter from the outer edge of the disk at $x_{\rm edge}=10^{3}$ with $\lambda_{\rm edge}=3.37$, $\mathcal{E}_{\rm edge}=4\times 10^{-4}$, and $\beta_{\rm edge}=10^{4}$, respectively. With this, we compute the global accretion solution that passes through the outer critical point at $x_{\rm out}=393.230$, connecting the outer edge ($x_{\rm edge}=10^{3}$) to the event horizon ($x_{\rm h}$) in the absence of thermal conduction ($\Upsilon_{\rm s}=0$). Thereafter, we calculate the energy ($\mathcal{E}$) profile of the inflow solution. The obtained results are shown in Fig. 6, where the variation of (a) Mach number ($M=v/C_{\rm s}$) and (b) local energy ($\mathcal{E}$) are plotted as function of radial coordinate ($x$). In both panels, the dotted (green) curve denotes the results for $\Upsilon_{\rm s}=0$. Next, we gradually increase $\Upsilon_{\rm s}$ while keeping the other inflow parameters fixed for a non-rotating black hole ($a_{\rm k}=0.0$) and obtain the global inflow solutions. The dashed (blue) and solid (red) curves represent results corresponding to $\Upsilon_{\rm s}=0.0025$, and $0.005$, respectively. The respective outer critical points are located at $x_{\rm out}=382.426$ and $371.241$, respectively. Evidently, thermal conduction influences the global accretion solutions, leading to an increase in the local energy of the flow.

### v2

@@ IV Astrophysical Implications
Further, X-ray luminosity ($L_{\rm x}$) for the same sample of AGNs is calculated in 0.2120.2-12 keV energy range from XMM-Newton data Rosen et al. (2016), which spans in the range of $10^{40.99}-10^{45.02}$ $\rm erg\,s^{-1}$. Using $L_{\rm x}$, we estimate the accretion power as $\dot{M}c^{2}=L_{\rm x}/\eta_{\rm acc}$, where $\eta_{\rm acc}$ denotes the accretion efficiency factor. Since the spin measurements of supermassive black holes (SMBHs) in these AGNs are not well constrained, we adopt a maximum efficiency of $\eta_{\rm acc}=0.3$ in this study Thorne (1974). Thereafter, we employ the accretion-ejection model formalism to calculate the maximum kinetic jet power Aktar et al. (2015); Nandi et al. (2018); Jana and Das (2024) using maximum mass outflow rate $R^{\rm max}_{\dot{m}}$ as
@@ V SUMMARY and CONCLUSION
In this study, we examine the effect of thermal conduction on the mass loss from a low angular momentum, advective, viscous, and magnetized accretion disk around a rotating black hole, for the first time to the best of our knowledge. In doing so, we assume the accreting plasma is collisionless and adopt the saturated form of thermal conduction Cowie and McKee (1977). In addition, we model the accretion disk as being threaded with a dominant toroidal magnetic field component Oda et al. (2007); Sarkar and Das (2016); Sarkar et al. (2018); Jana and Das (2024). To simplify the general relativistic complexities, we adopt a recently developed effective potential Dihingia et al. (2018) that satisfactorily mimics the space-time around a rotating black hole. Furthermore, we employ the relativistic equation of state (REoS) Chattopadhyay and Ryu (2009) to determine the thermodynamic quantities. With these assumptions, we solve the coupled governing equations for both inflow and outflow self-consistently in terms of the inflow parameters, namely energy, angular momentum, plasma-$\beta$, and the conduction parameter. We summarize the key findings of this study below:
@@
  • We compute the maximum mass outflow rate $R^{\rm max}_{\dot{\rm m}}$ as a function of $\Upsilon_{\rm s}$ for two extreme black hole spin parameters, $a_{\rm k}=0.0$ and $0.99$ (Fig. 4). Our results reveal that rapidly rotating black holes expel more mass from the magnetized disk in the form of outflows compared to the weakly rotating black holes. More precisely, when thermal conduction is active within the disk, $R_{\dot{\rm m}}^{\rm max}$ reaches up to $25\%$ for $a_{\rm k}=0.99$, whereas it remains below $22\%$ for $a_{\rm k}=0.0$ (see Fig. 4).
@@ Acknowledgements
$R_{\rm v}$ $=v+T_{\rm v},\quad R_{\Theta}=T_{\Theta},\quad R_{\lambda}=T_{\lambda},$
$R_{\beta}$ $=T_{\beta},\quad{\rm and}\quad R_{0}=T_{0}+2t\Theta/(\beta x)+d\Psi_{\rm e}^{\rm eff}/dx,$
@@
$E_{\rm v}$ $=t\Theta+\frac{5\Upsilon_{\rm s}(t\Theta)^{3/2}}{v},$
$\quad E_{\Theta}$ $=(vt/2+Nvt)-5\Upsilon_{\rm s}t^{3/2}\sqrt{\Theta},$
@@
$+\frac{5\Upsilon_{\rm s}(t\Theta)^{3/2}}{2}\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial\lambda},$
$E_{\beta}$ $=-\frac{vt\Theta}{2(1+\beta)\beta}-\frac{5\Upsilon_{\rm s}(t\Theta)^{3/2}}{2\beta(1+\beta)},$
@@
$+5\Upsilon_{\rm s}(t\Theta)^{3/2}\Big{(}-\frac{1}{x}+\frac{1}{2\Delta}\frac{d\Delta}{dx}+\frac{1}{2\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}\Big{)}.$
@@
$B_{0}$ $=\frac{1}{4}\Big{(}\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}-\frac{1}{\Delta}\frac{d\Delta}{dx})\Big{)}+\frac{1}{x}.$
@@
Figure 6: Variation of (a) Mach number ($M=v/C_{\rm s}$) and (b) energy ($\mathcal{E}$) of the inflowing matter with radial distance $(x)$ for $\Upsilon_{\rm s}$ = $0.0$ (green), $0.0025$ (blue) and $0.005$ (red). Here, we choose $a_{\rm k}=0$, $\alpha_{\rm B}=0.001$, and fix the input parameters at $x_{\rm edge}=1000$ as $\lambda_{\rm edge}=3.20$, $\mathcal{E}_{\rm edge}=2\times 10^{-4}$, and $\beta_{\rm edge}=10^{5}$. See the text for details.
Thermal conduction plays an important role in transferring heat energy within the accretion disk. Hence, it is worthy to examine the effect thermal conduction on the local energy of inflowing matter. In doing so, we choose $a_{\rm k}=0$, and $\alpha_{\rm B}=0.001$, and inject matter from the outer edge of the disk at $x_{\rm edge}=1000$ with $\lambda_{\rm edge}=3.20$, $\mathcal{E}_{\rm edge}=2\times 10^{-4}$, and $\beta_{\rm edge}=10^{5}$, respectively. With this, we compute the global accretion solution that passes through the outer critical point at $x_{\rm out}=800.166$, connecting the outer edge ($x_{\rm edge}=10^{3}$) to the event horizon ($x_{\rm h}$) in the absence of thermal conduction ($\Upsilon_{\rm s}=0$). Thereafter, we calculate the energy ($\mathcal{E}$) profile of the inflow solution. The obtained results are shown in Fig. 6, where the variation of (a) Mach number ($M=v/C_{\rm s}$) and (b) local energy ($\mathcal{E}$) are plotted as function of radial coordinate ($x$). In both panels, the dotted (green) curve denotes the results for $\Upsilon_{\rm s}=0$. Next, we gradually increase $\Upsilon_{\rm s}$ while keeping the other inflow parameters fixed for a non-rotating black hole ($a_{\rm k}=0.0$) and obtain the global inflow solutions. The dashed (blue) and solid (red) curves represent results corresponding to $\Upsilon_{\rm s}=0.0025$, and $0.005$, respectively. The respective outer critical points are located at $x_{\rm out}=788.559$ and $774.200$, respectively. Evidently, thermal conduction influences the global accretion solutions, leading to an increase in the local energy of the flow.